# PathVLM-R1: A Reinforcement Learning-Driven Reasoning Model for Pathology Visual-Language Tasks

### v1

@@ 1 Introduction
Traditional reinforcement learning methods, such as
Proximal Policy Optimization (PPO)[4]+Outcome Reward Model (ORM)[5], focus on the final outcome due to their black-box model structure, making it difficult to effectively verify the correctness of the reasoning process and thus falling short of the transparency and interpretability required for medical tasks. On the other hand, PPO[4]+Process Reward Model (PRM)[5] entails high annotation costs.
@@ 3.2 Redesigning Reward Mechanisms for Interpretability
However, test results based solely on format and accuracy rewards indicate that the quality of the generated reasoning chains is often suboptimal, frequently exhibiting logical leaps or misuse of medical arguments. Despite this, even low-quality reasoning chains can sometimes lead to correct results, which is detrimental to the model’s ability to find effective strategies. We realized that for open-ended tasks such as medical reasoning, accuracy rewards alone are insufficient. External models are needed to supervise the reasoning process. Given the scarcity of high-quality reasoning process data in the medical field, we referenced Reinforcement Learning with AI Feedback (RLAIF)[RLAIF, 27, 28] and DeepSeek-R1[25]’s training strategies to design a segmented approach.
@@ 4.2 Experimental Results
These findings underscore the importance of model parameterization and domain knowledge in R1 model training: models with larger parameters have higher potential for performance and improvement, and pre-training with domain knowledge through supervised fine-tuning significantly enhances data efficiency and optimization in subsequent RL steps. This is particularly evident in domains like medical scenarios, which differ substantially from common modalities and lack manually annotated data. Other reasoning models for medical imaging modalities, such as Med-R1[35] and MedVLM-R1[19], employ a rule-based reward function directly on the base model, optimized using the GRPO strategy. While this approach has also surpassed their foundational models and the supervised fine-tuning baseline, the results are suboptimal; MedVLM-R1 even failed in testing on pathology images. Our experimental results indicate that our post-training approach, which involves "pathological knowledge infusion and reasoning capability construction," can achieve superior performance and higher efficiency. Figure LABEL:fig:1 illustrates the reasoning process and final results for the same question before and after adding cross-modal procedural supervision rewards. PathVLM-R1 achieves better outcomes due to its superior reasoning capabilities.

### v2

@@ 1 Introduction
Figure 1: Comparison of the final PathVLM-R1 model and its variant PathVLM-$\beta$. For the same pathology image question-answering task, the PathVLM-R1, which incorporates cross-modal procedural rewards, demonstrates superior rigor in reasoning and accuracy in medical knowledge compared to the PathVLM-$\beta$, which utilizes only accuracy rewards and format rewards. Specifically, in this example, PathVLM-$\beta$ only focuses on differences in nuclear size while neglecting that "pleomorphism" includes multiple aspects of variation, such as nuclear shape and structure. In contrast, PathVLM integrates multiple nuclear features, accurately grasping the definition of "pleomorphism" with a more comprehensive thought process and stronger rigor.
@@
Traditional reinforcement learning methods, such as Proximal Policy Optimization (PPO)[4]+Outcome Reward Model (ORM)[5], focus on the final outcome due to their black-box model structure, making it difficult to effectively verify the correctness of the reasoning process and thus falling short of the transparency and interpretability required for medical tasks. On the other hand, PPO[4]+Process Reward Model (PRM)[5] entails high annotation costs.
@@ 3.2 Redesigning Reward Mechanisms for Interpretability
However, test results based solely on format and accuracy rewards indicate that the quality of the generated reasoning chains is often suboptimal, frequently exhibiting logical leaps or misuse of medical arguments. Despite this, even low-quality reasoning chains can sometimes lead to correct results, which is detrimental to the model’s ability to find effective strategies. We realized that for open-ended tasks such as medical reasoning, accuracy rewards alone are insufficient. External models are needed to supervise the reasoning process. Given the scarcity of high-quality reasoning process data in the medical field, we referenced Reinforcement Learning with AI Feedback (RLAIF)[26, 27, 28] and DeepSeek-R1[25]’s training strategies to design a segmented approach.
@@ 4.2 Experimental Results
These findings underscore the importance of model parameterization and domain knowledge in R1 model training: models with larger parameters have higher potential for performance and improvement, and pre-training with domain knowledge through supervised fine-tuning significantly enhances data efficiency and optimization in subsequent RL steps. This is particularly evident in domains like medical scenarios, which differ substantially from common modalities and lack manually annotated data. Other reasoning models for medical imaging modalities, such as Med-R1[35] and MedVLM-R1[19], employ a rule-based reward function directly on the base model, optimized using the GRPO strategy. While this approach has also surpassed their foundational models and the supervised fine-tuning baseline, the results are suboptimal; MedVLM-R1 even failed in testing on pathology images. Our experimental results indicate that our post-training approach, which involves "pathological knowledge infusion and reasoning capability construction," can achieve superior performance and higher efficiency. Figure 1 illustrates the reasoning process and final results for the same question before and after adding cross-modal procedural supervision rewards. PathVLM-R1 achieves better outcomes due to its superior reasoning capabilities.